OpenStudy (anonymous):

13/40+2/5+19/40 would best be solved by associative or distributive propert?

OpenStudy (compassionate):

\[\frac{ 13 }{ 40 } + \frac{ 2 }{ 5 } + \frac{ 19 }{ 40 }\] The associative property has to do with addition. It says x + y = y + x The distributive property has to do with multiplicaton. It says x(y + z) = xy + xz

OpenStudy (catlover5925):

associative, because you are adding, distributive would be used if you had something to distribute EX: 4(2x + 9)

OpenStudy (compassionate):

Cat, lets not give out answers. It hinders the learning process.

OpenStudy (catlover5925):

im so sorry I have a bad habbit of that when I don't know how to explain it

OpenStudy (anonymous):

can somebody please help me with my question?